Enough of a monoprotic acid is dissolved in water
Answer #1

given pH= 2.59

-log [H+] =2.59

[H+] =10-2.59 =0.00257

for mono protic acid   HA ---> H+ +A-

Ka= [H+] [A-]/[HA] Ka =equilirium constant

[H+] =[A-] = 0.00257 at equilibrium and [HA] at equulibrium= 0.0165-0.00257=0.01393

Ka= 0.00257*0.00257 0.01393=0.000474